### v3

# Yilun ${}^{\heartsuit}$, Xingyao Zhang, Arman ${}^{\heartsuit}$, Mark ${}^{\heartsuit}$

Abstract
Large language models (LLMs), despite their remarkable progress across various general domains, encounter significant barriers in medicine and healthcare. This field faces unique challenges such as domain-specific terminologies and reasoning over specialized knowledge. To address these issues, we propose a novel Multi-disciplinary Collaboration (MC) framework for the medical domain that leverages LLM-based agents in a role-playing setting that participate in a collaborative multi-round discussion, thereby enhancing LLM proficiency and reasoning capabilities. This training-free framework encompasses five critical steps: gathering domain experts, proposing individual analyses, summarising these analyses into a report, iterating over discussions until a consensus is reached, and ultimately making a decision. Our work focuses on the zero-shot setting, which is applicable in real-world scenarios. Experimental results on nine datasets (MedQA, MedMCQA, PubMedQA, and six subtasks from MMLU) establish that our proposed MC framework excels at mining and harnessing the medical expertise within LLMs, as well as extending its reasoning abilities. Our code can be found at https://github.com/gersteinlab/MedAgents.
[…]
MedAgents: Large Language Models as Collaborators for Zero-shot Medical Reasoning
Xiangru ${}^{\heartsuit}$thanks:  Equal contribution. , Anni ${}^{\spadesuit}$11footnotemark: 1 , Zhuosheng ${}^{\spadesuit}$, Ziming Li, Yilun ${}^{\heartsuit}$, Xingyao Zhang, Arman ${}^{\heartsuit}$, Mark ${}^{\heartsuit}$$^{\heartsuit}$ale University  ${}^{\spadesuit}$hanghai Jiao Tong University xiangru.tang@yale.edu
1 Introduction
Large language models (LLMs) (Brown et al., 2020; Scao et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023; OpenAI, 2023) have exhibited notable generalization abilities across a wide range of tasks and applications (Lu et al., 2023; Zhou et al., 2023; Park et al., 2023), with these capabilities stemming from their extensive training on vast comprehensive corpora covering diverse topics. However, in real-world scenarios, LLMs are inclined to encounter domain-specific tasks that necessitate a combination of domain expertise and complex reasoning abilities (Moor et al., 2023; Wu et al., 2023b; Singhal et al., 2023a; Yang et al., 2023). Amidst this backdrop, a noteworthy research topic lies in the adoption of LLMs in the medical field, which has gained increasing prominence recently (Zhang et al., 2023b; Bao et al., 2023; Singhal et al., 2023a).
Two major challenges prevent LLMs from effectively handling tasks in the medical sphere: (i) Limited volumn and specificity of medical training data compared to the vast general text data, due to cost and privacy concerns (Thirunavukarasu et al., 2023). [^1] Although Med-PaLM 2 (Singhal et al., 2023b) serves as a specialized medical LLM fine-tuned on the basis of PaLM 2, it is closed-sourced and not publicly accessible yet. (ii) The demand for extensive domain knowledge (Schmidt and Rikers, 2007) and advanced reasoning skills (Liévin et al., 2022) makes eliciting medical expertise via simple prompting challenging (Kung et al., 2023; Singhal et al., 2023a). Although numerous attempts, particularly within math and coding, have been made to enhance prompting methods (Besta et al., 2023; Lála et al., 2023; Wang et al., 2023b), strategies used in the medical field have been shown to induce hallucination (Umapathi et al., 2023; Harris, 2023; Ji et al., 2023), indicating the need for more robust approaches.
Figure 1: Diagram of our proposed MC framework. Given a medical question as input, the framework performs reasoning in five stages: (i) expert gathering, (ii) analysis proposition, (iii) report summarization, (iv) collaborative consultation, and (v) decision making.
Meanwhile, recent research has surprisingly witnessed the success of multi-agent collaboration (Xi et al., 2023; Wang et al., 2023d) which highlights the simulation of human activities (Du et al., 2023; Liang et al., 2023; Park et al., 2023) and optimizes the collective power of multiple agents (Chen et al., 2023; Li et al., 2023d; Hong et al., 2023). Through such design, the expertise implicitly embedded within LLMs, or that the model has encountered during its training, which may not be readily accessible via traditional prompting, is effectively brought to the fore. This process subsequently enhances the model’s reasoning capabilities throughout multiple rounds of interaction (Wang et al., 2023c; Fu et al., 2023).
Motivated by these notions, we pioneer a Multi-disciplinary Collaboration (MC) framework specifically tailored to the clinical domain. Our objective centers on unveiling the intrinsic medical knowledge embedded in LLMs and reinforcing reasoning proficiency in a training-free manner. As shown in Figure 1, the MC framework gathers experts from diverse disciplines and reaches consistent conclusions through collaborative discussions.
Based on our MC framework, we conduct experiments on nine datasets, including MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019) and six medical subtasks from MMLU (Hendrycks et al., 2020). [^2] We follow the evaluation setting from Faln-PaLM (Singhal et al., 2023a). To better align with real-world application scenarios, our study focuses on the zero-shot setting, which can serve as a plug-and-play method to supplement existing medical LLMs such as Med-PaLM 2 (Singhal et al., 2023b). Encouragingly, our proposed approach outperforms settings for both chain-of-thought (CoT) and self-consistency (SC) prompting methods. Most notably, our approach achieves better performance under the zero-shot setting compared with the 5-shot baselines.
Based on our results, we further investigate the influence of agent numbers and conduct human evaluations to pinpoint the limitations and issues prevalent in our approach. We find four common categories of errors: (i) lack of domain knowledge; (ii) mis-retrieval of domain knowledge; (iii) consistency errors; and (iv) CoT errors. Targeted refinements focused on mitigating these particular shortcomings would enhance the model’s proficiency and reliability.
Our contributions are summarized as follows:
(i) To the best of our knowledge, we are the first to propose a multi-agent framework within the medical domain.
(ii) Our proposed MC framework enjoys enhanced faithfulness and interpretability by harnessing role-playing and collaborative agent discussion. Examples illustrating interpretability are shown in Appendix B.
(iii) Experimental results on nine datasets demonstrate the general effectiveness of our proposed MC framework. Besides, we identify and categorize common error types in our approach through rigorous human evaluation to shed light on future studies.
2 Related Work
2.1 LLMs in Medical Domains
Recent years have seen remarkable progress in the application of LLMs (Wu et al., 2023b; Singhal et al., 2023a; Yang et al., 2023), with a particularly notable impact on the medical field (Bao et al., 2023; Nori et al., 2023; Rosoł et al., 2023). Although LLMs have demonstrated their potential in distinct medical applications encompassing diagnostics (Singhal et al., 2023a; Han et al., 2023), genetics (Duong and Solomon, 2023; Jin et al., 2023), pharmacist (Liu et al., 2023), and medical evidence summarization (Tang et al., 2023b, a; Shaib et al., 2023), concerns persist when LLMs encounter clinical inquiries that demand intricate medical expertise and decent reasoning abilities (Umapathi et al., 2023; Singhal et al., 2023a). Thus, it is of crucial importance to further arm LLMs with enhanced clinical reasoning capabilities. Currently, there are two major lines of research on LLMs in medical domains, tool-augmented methods and instruction-tuning methods.
For tool-augmented approaches, recent studies rely on external tools to acquire additional information for clinical reasoning. For instance, GeneGPT (Jin et al., 2023) guided LLMs to leverage the Web APIs of the National Center for Biotechnology Information (NCBI) to meet various biomedical information needs. Zakka et al. (2023) proposed Almanac, a framework that is augmented with retrieval capabilities for medical guidelines and treatment recommendations. Kang et al. (2023) introduced a method named KARD to improve small LMs on specific domain knowledge by fine-tuning small LMs on the rationales generated from LLMs and augmenting small LMs with external knowledge from a non-parametric memory.
Current instruction tuning research predominantly leverages external clinical knowledge bases and self-prompted data to obtain instruction datasets (Tu et al., 2023; Zhang et al., 2023a; Singhal et al., 2023b; Tang et al., 2023c). These datasets are then employed to fine-tune LLMs within the medical field (Singhal et al., 2023b). Some of these models utilize a wide array of datasets collected from medical and biomedical literature, fine-tuned with specialized or open-ended instruction data (Li et al., 2023a; Singhal et al., 2023b). Others focus on specific areas such as traditional Chinese medicine or large-scale, diverse medical instruction data to enhance their medical proficiency (Tan et al., 2023; Zhang et al., 2023b). Unlike these methods, our work emphasizes harnessing latent medical knowledge intrinsic to LLMs and improving reasoning in a training-free setting.
2.2 LLM-based Multi-agent Collaboration
The development of LLM-based agents has made significant progress in the community by endowing LLMs with the ability to perceive surroundings and make decisions individually (Wang et al., 2023a; Yao et al., 2022; Nakajima, 2023; Xie et al., 2023; Zhou et al., 2023). Beyond the initial single-agent mode, the multi-agent pattern has garnered increasing attention recently (Xi et al., 2023; Li et al., 2023d; Hong et al., 2023) which further explores the potential of LLM-based agents by learning from multi-turn feedback and cooperation. In essence, the key to LLM-based multi-agent collaboration is the simulation of human activities such as role-playing (Wang et al., 2023d; Hong et al., 2023) and communication (Wu et al., 2023a; Qian et al., 2023; Li et al., 2023b, c). For instance, Solo Performance Prompting (SPP) (Wang et al., 2023d) managed to combine the strengths of multiple minds to improve performance by dynamically identifying and engaging multiple personas throughout task-solving. Camel (Li et al., 2023b) leveraged role-playing to enable chat agents to communicate with each other for task completion. Several recent works attempt to incorporate adversarial collaboration including debates (Du et al., 2023; Xiong et al., 2023) and negotiation (Fu et al., 2023) among multiple agents to further boost performance. Liang et al. (2023) proposed a multi-agent debate framework in which various agents put forward their statements in a tit for tat pattern. Inspired by the multi-disciplinary consultation mechanism which is common and effective in hospitals, we are thus inspired to apply this mechanism to medical reasoning tasks through LLM-based multi-agent collaboration.
Figure 2: Illustrative example of our proposed multi-disciplinary collaboration framework.
3 Method
This section presents the details of our proposed Multi-disciplinary Collaboration (MC) framework. Figure 1 and 2 give an overview and an illustrative example of its pipeline, respectively. Our proposed MC framework works in five stages: (i) expert gathering: assemble experts from various disciplines based on the clinical question; (ii) analysis proposition: domain experts present their own analyses with their expertise; (iii) report summarization: develop a report summary on the basis of previous analyses; (iv) collaborative consultation: hold a consultation over the summarized report with the experts. The report will be revised repeatedly until every expert has given their approval. (v) decision making: derive a final decision from the unanimous report. [^3] Details about all guideline prompts and roles are shown in Section D for clarification.
3.1 Expert Gathering
Given a clinical question $q$ and a set of options $op=\left\{o_{1},o_{2},\ldots,o_{k}\right\}$ where $k$ is the number of options, the goal of the Expert Gathering stage is to recruit a group of question domain experts $\mathcal{QD}=\left\{qd_{1},qd_{2},\ldots,qd_{m}\right\}$ and option domain experts $\mathcal{OD}=\left\{od_{1},od_{2},\ldots,od_{n}\right\}$ where $m$ and $n$ represent the number of question domain experts and option domain experts, respectively. Specifically, we assign a role to the model and provide instructions to guide the model output to the corresponding domains based on the input question and options, respectively:
$\begin{split}\mathcal{QD}&=\text{LLM}\left(q,\texttt{r}_{\texttt{qd}},\texttt{prompt}_{\texttt{qd}}\right),\\ \mathcal{OD}&=\text{LLM}\left(q,op,\texttt{r}_{\texttt{od}},\texttt{prompt}_{\texttt{od}}\right),\\ \end{split}$ (1)
where $\left(\texttt{r}_{\texttt{qd}},\texttt{prompt}_{\texttt{qd}}\right)$ and $\left(\texttt{r}_{\texttt{od}},\texttt{prompt}_{\texttt{od}}\right)$ stand for the system role and guideline prompt to gather domain experts for the question $q$ and options $op$.
Dataset Format Choice Testing Size Domain
MedQA Question + Answer A/B/C/D 1273 US Medical Licensing Examination
MedMCQA Question + Answer A/B/C/D and Explanations 6.1K AIIMS and NEET PG entrance exams
PubMedQA Question + Context + Answer Yes/No/Maybe 500 PubMed paper abstracts
MMLU Question + Answer A/B/C/D 1089 Graduate Record Examination
& US Medical Licensing Examination
Table 1: Summary of the Datasets. Part of the values are from the appendix of Singhal et al. (2023a).
3.2 Analysis Proposition
After gathering domain experts for the question $q$ and options $op$, we aim to inquire experts to generate corresponding analyses prepared for later reasoning: $\mathcal{QA}=\{qa_{1},qa_{2},\ldots,qa_{m}\}$ and $\mathcal{OA}=\{oa_{1},oa_{2},\ldots,oa_{n}\}$.
Question Analyses
Given a question $q$ and a question domain $qd_{i}\in\mathcal{QD}$, we ask LLM to serve as an expert specialized in domain $qd_{i}$ and derive the analyses for the question $q$ following the guideline prompt $\texttt{prompt}_{\texttt{qa}}$:
$qa_{i}=\text{LLM}\left(q,qd_{i},\texttt{r}_{\texttt{qa}},\texttt{prompt}_{\texttt{qa}}\right).$ (2)
Option Analyses
Now that we have an option domain $od_{i}$ and question analyses $\mathcal{QA}$, we can further analyze the options by taking into account both the relationship between the options and the relationship between the options and question. Concretely, we deliver the question $q$, the options $op$, a specific option domain $od_{i}\in\mathcal{OD}$, and the question analyses $\mathcal{QA}$ to the LLM:
$oa_{i}=\text{LLM}\left(q,op,od_{i},\mathcal{QA},\texttt{r}_{\texttt{oa}},\texttt{prompt}_{\texttt{oa}}\right).$ (3)
3.3 Report Summarization
In the Report Summarization stage, we attempt to summarize and synthesize previous analyses from various domain experts $\mathcal{QA}\cup\mathcal{OA}$. Given question analyses $\mathcal{QA}$ and option analyses $\mathcal{OA}$, we ask LLMs to play the role of a medical report assistant, allowing it to generate a synthesized report by extracting key knowledge and total analysis based on previous analyses:
$Repo=\text{LLM}\left(\mathcal{QA},\mathcal{OA},\texttt{r}_{\texttt{rs}},\texttt{prompt}_{\texttt{rs}}\right).$ (4)
Input: Domain experts $D=\{d_{1},...,d_{n}\}$, initial report $R_{0}$, Model $\mathcal{M}$, maximum attempts $k$, prompts $\{p_{vote},p_{mod},p_{rev}\}$
Output: Final report $R_{f}$
// Initialize variables
$nocon\_flag\leftarrow True$, $n_{try}\leftarrow 0$
$R_{cur}\leftarrow R_{0}$, $Mods\leftarrow\emptyset$
// Iterative review
while $nocon\_flag$ is $True$ and $n_{try}<k$ do
       $n_{try}\leftarrow n_{try}+1$
       $nocon\_flag\leftarrow False$
       // vote for the report
       for $i$ in $1,...,n$ do
             $vote_{i}\leftarrow\mathcal{M}(R_{cur},d_{i},p_{vote})$
             // propose modifications
             if $vote_{i}$ is $no$ then
                   $Mod_{i}\leftarrow\mathcal{M}(R_{cur},d_{i},p_{mod})$
                   Update $Mods$ with $Mod_{i}$
                   $nocon\_flag\leftarrow True$
             end if
       end for
      // modify the report
       if $nocon\_flag$ is $True$ then
             $R_{cur}\leftarrow\mathcal{M}(R_{cur},Mods,p_{rev})$
       end if
end while
return $R_{f}\leftarrow R_{cur}$
Algorithm 1 Collaborative Consultation
3.4 Collaborative Consultation
Since we have a preliminary summary report $Repo$, the objective of the Collaborative Consultation stage is to engage distinct domain experts in multiple rounds of discussions and ultimately render a summary report that is recognized by all experts. The overall procedure of this phase is presented in Algorithm 1. During each round of discussions, the experts give their votes (yes/no) as well as modification opinions if they vote no for the current report. Afterward, the report will be revised based on the modification opinions. Specifically, during the $i$-th round of discussion, we note the modification comments from the experts as $Mod_{i}$, then we can acquire the updated report as $Repo_{i}=\text{LLM}\left(Repo_{i-1},Mod_{i},\texttt{prompt}_{\texttt{mod}}\right)$. In this way, the discussions are held iteratively until all experts vote yes for the final report $Repo_{f}$.
3.5 Decision Making
In the end, we demand LLM act as a medical decision maker to derive the final answer to the clinical question $q$ referring to the unanimous report $Repo_{f}$:
$ans=\text{LLM}\left(q,op,Repo_{f},\texttt{r}_{\texttt{dm}},\texttt{prompt}_{\texttt{dm}}\right).$ (5)
Method MedQA MedMCQA PubMedQA Anatomy Clinical College Medical Professional College Avg.
knowledge medicine genetics medicine biology
      Flan-Palm
Few-shot CoT 60.3 53.6 77.2 66.7 77.0 83.3 75.0 76.5 71.1 71.2
Few-shot CoT + SC 67.6 57.6 75.2 71.9 80.4 88.9 74.0 83.5 76.3 75.0
      GPT-3.5
*few-shot setting
Few-shot 54.7 56.7 67.6 65.9 71.3 59.0 72.0 75.7 73.6 66.3
Few-shot CoT 55.3 54.7 71.4 48.1 65.7 55.5 57.0 69.5 61.1 59.8
Few-shot CoT + SC 62.1 58.3 73.4 70.4 76.2 69.8 78.0 79.0 77.2 71.6
*zero-shot setting
Zero-shot 54.3 56.3 73.7 61.5 76.2 63.6 74.0 75.4 75.0 67.8
Zero-shot CoT 44.3 47.3 61.3 63.7 61.9 53.2 66.0 62.1 65.3 58.3
Zero-shot CoT + SC 61.3 52.5 75.7 71.1 75.1 68.8 76.0 82.3 75.7 70.9
MC framework (Ours) 64.1 59.3 72.9 65.2 77.7 69.8 79.0 82.1 78.5 72.1
      GPT-4
*few-shot setting
Few-shot 76.6 70.1 73.4 79.3 89.5 75.6 93.0 91.5 91.7 82.3
Few-shot CoT 73.3 63.2 74.9 75.6 89.9 61.0 79.0 79.8 63.2 73.3
Few-shot CoT + SC 82.9 73.1 75.6 80.7 90.0 88.2 90.0 95.2 93.0 85.4
*zero-shot setting
Zero-shot 73.0 69.0 76.2 78.5 83.3 75.6 90.0 90.0 90.0 80.6
Zero-shot CoT 61.8 69.0 71.0 82.1 85.2 80.8 92.0 93.5 91.7 80.8
Zero-shot CoT + SC 74.5 70.1 75.3 80.0 86.3 81.2 93.0 94.8 91.7 83.0
MC framework (Ours) 83.7 74.8 76.8 83.5 91.0 87.6 93.0 96.0 94.3 86.7
Table 2: Main results (Acc). SC denotes the self-consistency prompting method. Results in bold are the best performances.
4 Experiments
4.1 Setup
Tasks and Datasets.
We evaluate our MC framework on three benchmark datasets MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019), as well as six subtasks most relevant to the medical domain from MMLU datasets  (Hendrycks et al., 2020) including anatomy, clinical knowledge, college medicine, medical genetics, professional medicine, and college biology. Table 1 summarizes the data statistics. More information about the evaluated datasets is presented in Appendix C.
Implementation.
We utilize the popular and publicly available GPT-3.5-Turbo and GPT-4 (OpenAI, 2023) from Azure OpenAI Service. [^4] https://learn.microsoft.com/en-us/azure/ai-services/openai/ All experiments are conducted in the zero-shot setting. The temperature is set to 1.0 and top_p to 1.0 for all generations. The iteration number and temperature of SC are 5 and 0.7, respectively. The number $k$ of options is 4 except for PubMedQA (3). The numbers of domain experts for the question and options are set as: $m=5,n=2$ except for PubMedQA ($m=4,n=2$). We randomly sample 300 examples for each dataset and conduct experiments on them. Statistically, the cost of our method is \$1.41 for 100 QA examples (about ¢1.4 per question) and the inference time per example is about $40s$ [^5] We acknowledge that our proposed method requires more cost compared with CoT or direct prompting. However, our approach is relatively cost-effective and the improved performance potentially leads to better health outcomes.
Baselines.
We have utilized models that are readily accessible through public APIs with the following baselines:
[…]
$\bullet$ Setting w/ CoT: Zero-shot CoT (Kojima et al., 2022) directly incorporates the prompt Let’s think step by step after a question to facilitate inference. Few-shot CoT (Wei et al., 2022) adopts comparable methodologies to Few-shot but distinguishes itself by integrating rationales before deducing the answer.
$\bullet$ Setting w/ SC: SC (Wang et al., 2022) serves as an additional sampling method on Zero-shot CoT and Few-shot CoT, which yields the majority answer by sampling multiple chains.
4.2 Main Results
Table 2 presents the main results on the nine datasets, including MedQA, MedMCQA, PubMedQA, and six subtasks from MMLU. We compare our method with several baselines in both zero-shot and few-shot settings. Notably, our proposed MC framework outperforms the zero-shot baseline methods by a large margin, indicating the effectiveness of our MC framework in real-world application scenarios. Furthermore, our approach achieves comparable performance under the zero-shot setting compared with the strong baseline Few-shot CoT+SC.
Interestingly, the introduction of CoT occasionally leads to a surprising degradation in performance. [^6] A more detailed analysis of CoT’s impact is provided in Appendix A. We have found that reliance on CoT in isolation can inadvertently result in hallucinations—spurious outputs typically associated with the misapplication of medical terminologies. In contrast, our multi-agent role-playing methodology effectively mitigates these issues, thus underscoring its potential as a more robust approach in medically oriented LLM applications.
Table 3: Ablation study for different processes. Anal: Analysis proposition, Summ: Report summarization, Cons: Collaborative consultation.
Method Accuracy(%)
Direct Prompting 49.0
CoT Prompting 55.0
[…]
 + Anal 62.0($\uparrow 7.0$)
 + Anal & Summ 65.0($\uparrow 10.0$)
 + Anal & Summ & Cons 67.0($\uparrow 12.0$)
Dataset MedQA MedMCQA PubMedQA MMLU
#Question agents 5 5 4 5
#Option agents 2 2 2 2
Table 4: Optimal number of agents on MedQA, MedMCQA, PubMedQA, and MMLU.
Figure 3: Influence of the number of question and option agents on various datasets.
Figure 4: Examples of error cases from MedQA and MedMCQA datasets in four major categories including lack of domain knowledge, mis-retrieval of domain knowledge, consistency errors, and CoT errors.
Figure 5: Ratio of different categories in error cases.
Table 3: Ablation study for different processes. Anal: Analysis proposition, Summ: Report summarization, Cons: Collaborative consultation.